OpenStudy (solomonzelman):

f(x)=x^3+x+31 Slowly, posting bit by bit, please.

OpenStudy (anonymous):

Well? :p

OpenStudy (anonymous):

:'c

OpenStudy (anonymous):

you want to derived?

OpenStudy (solomonzelman):

\(\huge\color{blue}{Can~~~someone~~teach~~me?}\)

OpenStudy (austinl):

What do you need to do?

OpenStudy (solomonzelman):

I don't need to do anything with it as in school, but I want to know how to find the zeros of this function.

OpenStudy (austinl):

Try to get it in this form. \(x(x\pm\underline{~~~})(x\pm\underline{~~~})\)

OpenStudy (solomonzelman):

How?

OpenStudy (anonymous):

if you substitute 0 to x F(x)=31?

OpenStudy (solomonzelman):

k, then....

OpenStudy (anonymous):

or \[F \prime (x) =3x^{2}+x \] then \[3x^{2}+x=0 \[x(3x+1)=0\] \[x=0 \] and \[x=-\frac{ 1 }{ 3 }\]

OpenStudy (solomonzelman):

what about 31, where did it go?

OpenStudy (anonymous):

31 when we don't derived F(x)

OpenStudy (anonymous):

That makes no sense. Sorry Solomon, I know you're upset with me. I missunderstood how this was going to work in the beginning. I apologize. But where did you get this problem?

OpenStudy (solomonzelman):

batman, did you just say that the question makes no sense? ganeshie8 gave it to me, as an example.

OpenStudy (anonymous):

No, I was referring to her explanation. It threw me off. I was just wondering, because it didn't even start off in the correct form: ax^2+bx+c. It cannot be factored. I tried to do the opposite of b thing, and that doesn't work either, well, it might, but it's a really awkward example. You should have started off with something like x^2+6x+8 or something..

OpenStudy (solomonzelman):

NO, I am trying to learn cubic functions, I know how to do things like ax^3+bx^2+cx+d, but this one is just ax^3+cx+d

OpenStudy (anonymous):

Well then. When we did stuff like that, you had to plug in a false number to replace that empty spot, because it has a peice missing. But this one is still a really hard example. Like, for yours, we would do x^3+0x^2+x+31. But normally, you would factor by grouping.

OpenStudy (solomonzelman):

but what do I do after plugging 0x^2 in?

OpenStudy (anonymous):

Like I said, normally we would factor by grouping... But this problem is really awkward.

OpenStudy (solomonzelman):

So how would be do this one, do you know?

OpenStudy (anonymous):

I'm looking it up. There are a few things I've found. There's one where you factor c, and plug in the factor's for all the x's. There's another where you do: \[\frac{ -b \pm \sqrt{b^{2}-3ac} }{ 3a }\] But I wish you could have started off with an easier example.

OpenStudy (solomonzelman):

I know how to do the easier once, hold on....

OpenStudy (anonymous):

Which way do you do the simpler ones?

OpenStudy (solomonzelman):

I find the zeros, as +- factors of the constant. So I would say that they are +-1, +-31, but none of them work.

OpenStudy (anonymous):

I am now fully determined to figure this out.

OpenStudy (solomonzelman):

Can someone just show me, without giving me mysterious tips?

OpenStudy (anonymous):

I just asked my goto, and he said, " I don't think it can be solved/factored. The roots aren't "real" There IS a quadratic formula equivalent for cubics but I think it's a lot longer and impractical by hand.

OpenStudy (anonymous):

Like I said, bad example. Why doesn't the guy who gae the problem to you show you how to do it?

OpenStudy (solomonzelman):

Whatever the roots are...

OpenStudy (solomonzelman):

Imaginary, complex, can you show me how to find them?

OpenStudy (anonymous):

*gave. Find what?

OpenStudy (solomonzelman):

\[f(x)=x^3+x+31~~~~~find~~the~~roots.\]